OpenStudy (anonymous):

1000 raffle tickets are sold for $1 each. One prize of $800 is to be awarded. a) Rena Condos purchases one ticket. Determine her expected value and b) Determine the fair price of a ticket.

OpenStudy (anonymous):

this what I got but I don't know if its correct for a) 800x(1/1000)+(-1)(999/1000) =800/1000-999/1000 =800-999/1000 =-199/1000=-0.199